OpenStudy (anonymous):

What is the slope between the two points (8, 15) and (3, 5)?

OpenStudy (anonymous):

General linear equation y=mx+c m=gradient c=y-intercept 1) First use the gradient formula m=y2-y1 / x2-x1 m= 5-15 / 3-8 m = .......... (you can do this) now use one point, lets choose (3,5) and the gradient we found m=.... and substitute back into the general equation y=mx + c we then are able to find c. Now we rewrite the whole equation with the gradient and c. Try it yourself and I will check it.

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

nope.

OpenStudy (anonymous):

(8,15) and (3,5) x1,y1 x2,y2 m(gradient)= 5-15 / 3-8 =........

OpenStudy (anonymous):

2

OpenStudy (anonymous):

yes that is correct.